# MOLLM: Multi-Objective Large Language Model for Molecular Design – Optimizing with Experts

### v1

@@ 4 Experiment
4.1 Task
The initial population plays a critical role in determining the final outcomes of genetic-based algorithms under a fixed computational budget. However, most prior studies have overlooked its significance. In practical applications, researchers often initialize the search with the best available molecules rather than selecting them entirely at random. To ensure a comprehensive and fair evaluation of our model, we conduct experiments on three distinct initialization scenarios: the top 100, bottom 100, and randomly sampled molecules from the ZINC 250K dataset, using their F-values as indicators. The best-initialization scenario assesses the model’s upper performance limit, the random initialization reflects common real-world usage, and the worst-initialization scenario presents a more challenging optimization task. We adhere to the PMO benchmark and operate within a budget of 5,000 oracle calls. For molecular property optimization, we focus on the following objectives: QED (drug-likeness), SA (synthetic accessibility), LogP (octanol-water partition coefficient), DRD2 (dopamine receptor D2 affinity), LogS (log of solubility), reduction potential, JNK3 (c-Jun N-terminal Kinase 3), and GSK3 (Glycogen Synthase Kinase 3 Beta). In addition to these well-defined objectives, we also include BBBP (Blood-Brain Barrier Permeability), a more complex and less predictable property influenced by multiple biological factors.
@@ 4.2 Metrics
Top 1 F & Mean Top 10 F: F (fitness) is the sum of the normalized objective values, which gives the direct representation of the strength of a molecule (Wang et al., 2024b). The weight in our experiment to each objective is the same.
(1) where is a molecule in SMILES form, k is the number of objectives, and is the weight and normalized objective value. If an objective is to be minimized, it will be transformed by . We give an equal weight to each objective.
Uniqueness: the fraction of valid generated molecules that are unique. A low uniqueness highlights repetitive molecule generation and a low level of distribution learning by the model (Bagal et al., 2021), while a high uniqueness value means that the model effectively explores novel molecules, the equation is blow:
(2) where is the number of repeated molecules, and is the total number of molecules proposed in history.
Validity: the fraction of molecules generated that are valid, it measures how well the model has learned the SMILES grammar and the valency of atoms (Bagal et al., 2021). The equation of validity is below:
(3) where is the number of valid molecules.
Structural Diversity: Structural diversity reflects the chemical diversity of the Pareto set and is computed by taking the average pairwise Tanimoto distance between Morgan fingerprints of molecules in the set (Benhenda, 2017). The equation of computing a set of molecules is:
(4) where is the set of molecules and is the tonimoto distance.
Efficiency: Efficiency is compared by the running time in hours, as well as LLM calls if application. It is a important metric when using LLM for inference, because querying LLM incurs high computational costs.
@@ 4.3 Baselines
To demonstrate the superiority and for fair comparison extensively, we choose SOTA models from a series of algorithms including GA, BO, MCMC, RL, DL and LLM-based method as our baselines. These algorithms are GB-GA, GB-BO, JT-VAE, MARS, REINVENT, MOLLEO, and recently proposed DyMol and Genetic-GFN which have achieved SOTA performance. More details and hyperparemeters of each baseline are provided in Appendix LABEL:apd:baseline. For a fair comparison, we use Chatgpt 4o for both MOLLM and MOLLEO. We use the default hyperparameters for GB-GA, JT-VAE, GB-BO, MARS, REINVENT defined in PMO benchmark (Gao et al., 2022). In terms of MOLLEO, DyMol and Genetic-GFN, we also use the default hyperparameters defined in their codes and papers. For fair comparison, the normalized objectives are applied for all methods, which also includes the correct optimizing direction.
@@ 4.4 Main Experiment Results
Top1 F | 4.048 | 3.817 | 3.665 | 3.907 | - | 4.096 | - | - | 4.187 |
Top10 F | 4.019 | 3.782 | 3.637 | 3.853 | - | 4.044 | - | - | 4.152 |
Uniqueness | 0.786 | 1.000 | 1.000 | 0.488 | - | 0.672 | - | - | 0.937 |
Validity | 1.000 | 1.000 | 1.000 | 1.000 | - | 0.930 | - | - | 0.915 |
Diversity | 0.583 | 0.847 | 1.000 | 0.826 | - | 0.656 | - | - | 0.556 |
@@
Top1 F | 3.941 | 3.923 | 4.015 | 3.924 | 4.092 | 4.098 | 4.232 | 4.157 | 4.276 |
Top10 F | 3.926 | 3.851 | 3.937 | 3.875 | 4.023 | 4.065 | 4.164 | 4.087 | 4.245 |
Uniqueness | 0.821 | 0.956 | 1.000 | 0.477 | 0.690 | 0.575 | 0.986 | 0.349 | 0.949 |
Validity | 1.000 | 1.000 | 1.000 | 0.999 | 0.979 | 0.938 | 1.000 | 0.998 | 0.900 |
Diversity | 0.623 | 0.778 | 0.717 | 0.819 | 0.640 | 0.570 | 0.581 | 0.653 | 0.529 |
@@
Top1 F | 0.948 0.941 | 1.901 1.887 | 2.901 2.891 | 3.901 3.890 | 4.276 4.098 | 5.183 4.964 |
Top10 F | 0.948 0.936 | 1.901 1.882 | 2.901 2.886 | 3.901 3.887 | 4.245 4.065 | 5.164 4.948 |
Uniqueness | 0.929 0.150 | 0.666 0.231 | 0.778 0.273 | 0.807 0.387 | 0.949 0.575 | 0.957 0.591 |
Validity | 0.796 0.159 | 0.962 0.552 | 0.946 0.803 | 0.946 0.783 | 0.900 0.938 | 0.890 0.926 |
Diversity | 0.538 0.865 | 0.450 0.646 | 0.510 0.627 | 0.375 0.614 | 0.529 0.573 | 0.529 0.611 |
@@
Following the experimental settings of MOLLEO (Wang et al., 2024b), we first conduct experiments to optimize five molecular properties simultaneously using molecules sampled from the ZINC 250K dataset. Among these objectives, three are minimized: SA, DRD2, and GSK3, while two are maximized: QED and JNK3. Each model is run with five different random seeds, and the final results are reported as the average over these runs. Since the initial population for REINVENT, DyMol, and Genetic-GFN cannot be explicitly set, these models are only evaluated in the randomly initialized scenario. The key evaluation metrics are top-1 fitness and average top-10 fitness, both of which directly reflect the sum of the normalized property values. To enhance clarity, the highest values in each metric are highlighted in Table 1. Our model demonstrates a significant improvement over other SOTA models across all three initialization cases, with a clear performance gap compared to the second-best approach. Notably, in both the worst-initialization and random-initialization scenarios, the mean top-10 F-value exceeds the top-1 F-value of the second-best model, highlighting the superior performance and convergence capabilities of our approach.
@@
In the best-initialization scenario, while the top-1 fitness of MOLLEO matches that of MOLLM, the mean top-10 fitness of MOLLM is noticeably higher than both the top-1 and mean top-10 fitness of all other models. Furthermore, our model maintains a uniqueness rate above 90%, whereas MOLLEO, despite being another LLM-based method, exhibits significantly lower uniqueness. This underscores the strong capability of MOLLM in effectively exploring the chemical space. The validity of generated molecules is also comparable to other models. Although our model exhibits relatively lower diversity among the top-100 molecules, we observe that models with higher diversity often achieve lower top fitness values. This suggests that direct comparisons of diversity may be less meaningful in this context but highlight a potential direction for future improvements. Across all three initialization settings, MOLLM consistently maintains higher diversity while achieving superior fitness values, demonstrating its robustness in molecular optimization.
5 Ablation Study
@@
In addition to the SOTA results from our main experiments involving the optimization of five objectives, we conduct further experiments with one to six objectives to assess the efficacy of MOLLM across varying optimization complexities and less predictable properties. Following this, we present an analysis of an interesting finding related to the experience pool utilized in our algorithm. Finally, we evaluate the impact of hyperparameters and demonstrate the effectiveness of our proposed components.
@@ 5.1 More Objectives
To further investigate MOLLM’s performance across different objective configurations, we conduct experiments using random initialization across scenarios with one to six objectives. The specific objective combinations are as follows:
1.
QED
2.
QED + SA
3.
QED + SA + DRD2
4.
QED + SA + DRD2 + GSK3
5.
QED + SA + DRD2 + GSK3 + JNK3
6.
QED + SA + DRD2 + GSK3 + JNK3 + BBBP
As the number of objectives increases, the performance gap between MOLLM and MOLLEO widens, particularly when optimizing more than four objectives, highlighting the superior capability of MOLLM in handling MOO. Additionally, MOLLM consistently achieves higher uniqueness and competitive validity compared to MOLLEO, while in MOLLEO these metrics tend to degrade significantly when optimizing fewer objectives. The consistently high uniqueness across all cases underscores the stability and effectiveness of MOLLM in optimization tasks with varying numbers of objectives. To further assess the robustness of MOLLM, we introduce BBBP (Blood-Brain Barrier Permeability) as a sixth objective, as it is a more complex and less predictable property with limited domain knowledge. Notably, despite the increased difficulty, MOLLM successfully generates a top 100 molecule set where all molecules are Blood-Brain Barrier Permeable (BBB+), demonstrating its strong adaptability and effectiveness in optimizing challenging molecular properties.
Method | LLM calls | Runing Time (hours) |
MOLLEO | 8517 | 7.32 |
MOLLM | 2908 | 0.52 |
Apart from that, without early stopping, MOLLM only uses nearly LLM calls compared to MOLLEO, more than even 14x faster than MOLLEO in run time to achieve significantly better results, as shown in Table 3.

### v2

@@ 4 Experiment
4.1 Task settings
The formulation of MOO using GA has often lacked rigor in previous studies. First, oracle calls should be restricted to ensure fair comparisons and practical applicability, since the evaluation of certain molecular properties requires costly experiments or specifically trained models, as noted in the Practical Molecular Optimization (PMO) benchmark Gao et al. (2022). Additionally, the initial population significantly impacts final performance under a fixed number of oracle calls, yet this factor has been largely overlooked in methods that incorporate genetic algorithms, such as MARS, MOLLEO, and GB-GA. In practical applications, researchers often initialize the search with the best available molecules rather than selecting them entirely at random. To ensure a comprehensive and fair evaluation of our model, we conduct experiments on three distinct initialization scenarios: the top 100, bottom 100, and randomly sampled molecules from the ZINC 250K dataset, using their F-values as indicators. The best-initialization scenario assesses the model’s upper performance limit, the random initialization reflects common real-world usage, and the worst-initialization scenario presents a more challenging optimization task. We adhere to the PMO benchmark and operate within a budget of 5,000 oracle calls. For molecular property optimization, we adopt the multi-objective task from MOLLEO: maximizing QED (drug-likeness) and JNK3 (c-Jun N-terminal Kinase 3), and minimizing SA (synthetic accessibility), LogP (octanol-water partition coefficient), DRD2 (dopamine receptor D2 affinity), LogS (log of solubility), reduction potential, and GSK3 (Glycogen Synthase Kinase 3 Beta). In addition to these well-defined objectives, we also include maximizing BBBP (Blood-Brain Barrier Permeability), a more complex and less predictable property influenced by multiple biological factors. In addition, we evaluated our method on the PMO benchmark, which includes structure-based optimization, name-based optimization, and property optimization. This benchmark features both single-objective tasks and multi-objective scenarios, such as the multi-parameter optimization of amlodipine, modeled after Amlodipine, a calcium channel blocker. This example task focuses on optimizing QED, logP, HBA/HBD, and SA, aiming to generate compounds that are realistic drug candidates while structurally distinct from the original molecule.
@@ 4.2 Metrics
Top 1 F & Mean Top 10 F: F (fitness) is the sum of the normalized objective values, which gives the direct representation of the strength of a molecule Wang et al. (2024b). The weight in our experiment to each objective is the same.
where is a molecule in SMILES form, k is the number of objectives, and is the weight and normalized objective value. If an objective is to be minimized, it will be transformed by . We give an equal weight to each objective. Uniqueness: the fraction of valid generated molecules that are unique. A low uniqueness highlights repetitive molecule generation and a low level of distribution learning by the model Bagal et al. (2021), while a high uniqueness value means that the model effectively explores novel molecules. Validity: the fraction of molecules generated that are valid, it measures how well the model has learned the SMILES grammar and the valency of atoms Bagal et al. (2021). Structural Diversity: Structural diversity reflects the chemical diversity of the Pareto set and is computed by taking the average pairwise Tanimoto distance between Morgan fingerprints of molecules in the set Benhenda (2017). Efficiency: Efficiency is compared by the running time in hours, as well as LLM calls if application. It is a important metric when using LLM for inference, because querying LLM incurs high computational costs.
@@ 4.3 Baselines
To demonstrate the superiority and for fair comparison extensively, we choose SOTA models from a series of algorithms including GA, BO, MCMC, RL, DL and LLM-based method as our baselines. These algorithms are GB-GA, GB-BO, JT-VAE, MARS, REINVENT, MOLLEO, and recently proposed DyMol and Genetic-GFN which have achieved SOTA performance. For a fair comparison, we use Chatgpt 4o for both MOLLM and MOLLEO. We use the default hyperparameters for GB-GA, JT-VAE, GB-BO, MARS, REINVENT defined in PMO benchmark Gao et al. (2022). In terms of MOLLEO, DyMol and Genetic-GFN, we also use the default hyperparameters defined in their codes and papers. For fair comparison, the normalized objectives are applied for all methods, which also includes the correct optimizing direction. Each model is run for 5 seeds independently and takes the average. On the PMO benchmark, we use the scores evaluated in the MOLLEO Wang et al. (2024b) paper, including the Augmented memory. We also notice the GuacaMol Brown et al. (2019) benchmark, but it does not constrain the oracle calls, which is unsuitable for our experiments.
@@ 4.4 Main Experiment Results
Top1 F | 4.048 | 3.838 | 3.665 | 3.891 | - | 4.096 | - | - | 4.229 |
Top10 F | 4.019 | 3.784 | 3.647 | 3.852 | - | 4.044 | - | - | 4.186 |
AUC-Top10 | 3.789 | 3.712 | 3.489 | 3.740 | - | 3.825 | - | - | 3.915 |
Hypervolume | 0.474 | 0.364 | 0.167 | 0.488 | - | 0.720 | - | - | 0.737 |
Uniqueness | 0.786 | 1.000 | 1.000 | 0.488 | - | 0.672 | - | - | 0.829 |
Validity | 1.000 | 1.000 | 1.000 | 1.000 | - | 0.930 | - | - | 0.937 |
Diversity | 0.583 | 0.847 | 1.000 | 0.826 | - | 0.656 | - | - | 0.603 |
@@
Top1 F | 4.017 | 3.874 | 4.003 | 3.931 | 4.230 | 4.190 | 4.232 | 4.299 | 4.413 |
Top10 F | 3.975 | 3.831 | 3.968 | 3.868 | 4.136 | 4.076 | 4.164 | 4.262 | 4.388 |
AUC-Top10 | 3.861 | 3.771 | 3.802 | 3.789 | 3.930 | 3.949 | 4.001 | 4.098 | 4.116 |
Hypervolume | 0.643 | 0.428 | 0.507 | 0.409 | 0.742 | 0.860 | 0.868 | 0.871 | 0.905 |
Uniqueness | 0.821 | 0.956 | 1.000 | 0.477 | 0.690 | 0.575 | 0.986 | 0.349 | 0.872 |
Validity | 1.000 | 1.000 | 1.000 | 0.999 | 0.979 | 0.938 | 1.000 | 0.998 | 0.908 |
Diversity | 0.623 | 0.778 | 0.717 | 0.819 | 0.640 | 0.570 | 0.581 | 0.633 | 0.441 |
@@
Top1 F | 0.948 0.941 | 1.901 1.887 | 2.901 2.891 | 3.901 3.890 | 4.413 4.098 | 5.183 4.964 |
Top10 F | 0.948 0.936 | 1.901 1.882 | 2.901 2.886 | 3.901 3.887 | 4.388 4.065 | 5.164 4.948 |
Uniqueness | 0.929 0.150 | 0.666 0.231 | 0.778 0.273 | 0.807 0.387 | 0.872 0.575 | 0.957 0.591 |
Validity | 0.796 0.159 | 0.962 0.552 | 0.946 0.803 | 0.946 0.783 | 0.908 0.938 | 0.890 0.926 |
Diversity | 0.538 0.865 | 0.450 0.646 | 0.510 0.627 | 0.375 0.614 | 0.441 0.573 | 0.529 0.611 |
@@
Following the experimental settings of MOLLEO Wang et al. (2024b), we first conduct experiments to optimize five molecular properties simultaneously using molecules sampled from the ZINC 250K dataset. Among these objectives, three are minimized: SA, DRD2, and GSK3, while two are maximized: QED and JNK3. Each model is run with five different random seeds, and the final results are reported as the average over these runs. Since the initial population for REINVENT, DyMol, and Genetic-GFN cannot be explicitly set, these models are only evaluated in the randomly initialized scenario. The key evaluation metrics are top-1 fitness and average top-10 fitness, both of which directly reflect the sum of the normalized property values. To enhance clarity, the highest values in each metric are highlighted in Table 1. Our model demonstrates a significant improvement over other SOTA models across all three initialization cases, with a clear performance gap compared to the second-best approach. In both random-initialization and worst-initialization scenarios, even the mean top-10 F-value of MOLLM exceeds the top-1 F-value of the second-best model, highlighting the superior performance and convergence capabilities of our approach. We evaluate the novelty of our output molecules to ZINC 250K dataset. Notably, every run for all scenarios outputs 100% novel molecules that do not exist in ZINC 250K.
@@
In the best-initialization scenario, while the top-1 fitness of MOLLEO matches that of MOLLM, the mean top-10 fitness of MOLLM is noticeably higher than both the top-1 and mean top-10 fitness of all other models. Furthermore, our model maintains a uniqueness rate above 90%, whereas MOLLEO, despite being another LLM-based method, exhibits significantly lower uniqueness. This underscores the strong capability of MOLLM in effectively exploring the chemical space. The validity of generated molecules is also comparable to other models. Although our model exhibits relatively lower diversity among the top-100 molecules, we observe that models with higher diversity often achieve lower top fitness values. This suggests that direct comparisons of diversity is less meaningful in this context but highlight a potential direction for future improvements.
4.5 Property values of 5-objective optimization
Method | SA | DRD2 | GSK3 | QED | JNK3 |
Starting population | |||||
(random sampled) | 3.021 | 0.009 | 0.026 | 0.730 | 0.015 |
MOLLM | 1.990 | 0.001 | 0.069 | 0.828 | 0.646 |
MOLLEO | 1.861 | 0.001 | 0.013 | 0.919 | 0.133 |
Genetic-GFN | 1.998 | 0.001 | 0.018 | 0.865 | 0.408 |
DyMol | 2.125 | 0.001 | 0.019 | 0.763 | 0.409 |
REINVENT | 1.878 | 0.001 | 0.012 | 0.900 | 0.235 |
MARS | 2.046 | 0.002 | 0.010 | 0.862 | 0.010 |
GB-BO | 1.725 | 0.001 | 0.018 | 0.887 | 0.111 |
JT-VAE | 2.144 | 0.001 | 0.008 | 0.904 | 0.018 |
GB-GA | 1.670 | 0.001 | 0.036 | 0.901 | 0.107 |
We present the objective values for the five-objective optimization task in Table 10. Compared to the initial population by random sampling, all models achieve notable improvements across nearly all objectives. Notably, while MOLLM achieves competitive performance on SA, DRD2, GSK3, and QED, it demonstrates a significant advantage in optimizing JNK3, outperforming the second-best model by a substantial margin. This result highlights MOLLM’s ability to effectively balance conflicting objectives, as improving JNK3 alone is relatively straightforward, but simultaneously enhancing JNK3 while reducing DRD2 and GSK3 remains a much more challenging task, which the other models fail to address adequately.
4.6 Results on PMO
Task type | Objective() | REINVENT | Augmented Memory | Graph GA | GP BO | MOLLEO | Genetic GFN | MOLLM (Ours) |
Property optimization | QED | 0.941 | 0.941 | 0.940 | 0.937 | 0.948 | 0.942 | 0.943 |
JNK3 | 0.783 | 0.773 | 0.553 | 0.564 | 0.790 | 0.764 | 0.822 | |
DRD2 | 0.945 | 0.962 | 0.964 | 0.923 | 0.968 | 0.974 | 0.974 | |
GSK3 | 0.865 | 0.889 | 0.788 | 0.851 | 0.863 | 0.881 | 0.836 | |
Name-based optimization | mestranol_similarity | 0.618 | 0.764 | 0.579 | 0.627 | 0.972 | 0.708 | 0.904 |
albuterol_similarity | 0.896 | 0.918 | 0.874 | 0.902 | 0.985 | 0.949 | 0.985 | |
thiothixene_rediscovery | 0.534 | 0.562 | 0.479 | 0.559 | 0.727 | 0.583 | 0.547 | |
celecoxib_rediscovery | 0.716 | 0.784 | 0.582 | 0.728 | 0.864 | 0.802 | 0.869 | |
troglitazone_rediscovery | 0.452 | 0.556 | 0.377 | 0.405 | 0.562 | 0.511 | 0.707 | |
perindopril_mpo | 0.537 | 0.598 | 0.538 | 0.493 | 0.600 | 0.595 | 0.646 | |
ranolazine_mpo | 0.760 | 0.802 | 0.728 | 0.735 | 0.769 | 0.819 | 0.843 | |
sitagliptin_mpo | 0.021 | 0.479 | 0.433 | 0.186 | 0.584 | 0.634 | 0.587 | |
amlodipine_mpo | 0.642 | 0.686 | 0.625 | 0.552 | 0.773 | 0.761 | 0.704 | |
fexofenadine_mpo | 0.769 | 0.686 | 0.779 | 0.745 | 0.847 | 0.856 | 0.856 | |
osimertinib_mpo | 0.834 | 0.856 | 0.808 | 0.762 | 0.835 | 0.860 | 0.888 | |
zaleplon_mpo | 0.347 | 0.438 | 0.456 | 0.272 | 0.510 | 0.552 | 0.571 | |
median1 | 0.372 | 0.335 | 0.287 | 0.325 | 0.352 | 0.379 | 0.339 | |
median2 | 0.294 | 0.290 | 0.229 | 0.308 | 0.275 | 0.294 | 0.310 | |
Structure-based optimization | isomers_c7h8n2o2 | 0.842 | 0.954 | 0.949 | 0.662 | 0.984 | 0.969 | 0.977 |
isomers_c9h10n2o2pf2cl | 0.642 | 0.830 | 0.719 | 0.469 | 0.874 | 0.897 | 0.893 | |
deco_hop | 0.666 | 0.688 | 0.619 | 0.629 | 0.942 | 0.733 | 0.938 | |
scaffold_hop | 0.560 | 0.565 | 0.517 | 0.548 | 0.971 | 0.615 | 0.967 | |
valsartan_smarts | 0.000 | 0.000 | 0.000 | 0.000 | 0.867 | 0.135 | 0.904 | |
Total () | 14.036 | 15.356 | 13.823 | 13.182 | 17.862 | 16.213 | 18.010 | |
Rank () | 5 | 4 | 6 | 7 | 2 | 3 | 1 |
On the PMO benchmark, we comprehensively evaluate our model across both single-objective and multi-objective optimization tasks, including property, name-based, and structure-based optimization, as shown in Table 4. The best values are highlighted in bold, with the next top three values underlined. Our model ranks first in 12 out of 23 tasks and achieves the highest total score and overall ranking, demonstrating its superiority across diverse objectives and consistent SOTA performance.
5 Empirical Study
@@
In addition to the SOTA results from our main experiments involving the optimization of five objectives, we conduct further experiments with one to six objectives to assess the efficacy of MOLLM across varying optimization complexities and less predictable properties. Following this, we present an analysis of an interesting finding related to the experience pool utilized in our algorithm. We also conducted experiments of the effects of different output molecules per query in Section 7.3 in the Appendix.
@@ 5.1 More Objectives
To further investigate MOLLM’s performance across different objective configurations, we conduct experiments using random initialization across scenarios with one to six objectives, starting with only QED to QED + SA + DRD2 + GSK3 + JNK3 + BBBP. As the number of objectives increases, the performance gap between MOLLM and MOLLEO widens, particularly when optimizing more than four objectives, highlighting the superior capability of MOLLM in handling MOO. Additionally, MOLLM consistently achieves higher uniqueness and competitive validity compared to MOLLEO, while in MOLLEO these metrics tend to degrade significantly when optimizing fewer objectives. The consistently high uniqueness across all cases underscores the stability and effectiveness of MOLLM in optimization tasks with varying numbers of objectives. To further assess the robustness of MOLLM, we introduce BBBP (Blood-Brain Barrier Permeability) as a sixth objective, as it is a more complex and less predictable property with limited domain knowledge. Notably, despite the increased difficulty, MOLLM successfully generates a top 100 molecule set where all molecules are Blood-Brain Barrier Permeable (BBB+), demonstrating its strong adaptability and effectiveness in optimizing challenging molecular properties.
Method | MOLLM | MOLLEO | Graph GA | Gp-BO | Genetic-GFN | MARS | JT-VAE | DyMol | REINVENT |
Running time(hours) | 0.52 | 7.32 | 1.04 | 0.68 | 0.45 | 0.69 | 3.08 | 0.28 | 0.52 |
LLM calls | 2908 | 8517 | - | - | - | - | - | - | - |
Apart from that, without early stopping, MOLLM only uses nearly LLM calls compared to MOLLEO, more than even 14x faster than MOLLEO in run time to achieve significantly better results, as shown in Table 5. The running time is also competitive to other methods that are not based on LLMs.